# MLR-Bench: Evaluating AI Agents on Open-Ended Machine Learning Research

### v1

@@ 1 Introduction
  • Comprehensive Evaluation and Insights: We apply MLR-Bench to evaluate 6 frontier models, providing comparative insights into their strengths and limitations in scientific discovery. Our analysis highlights critical failure modes, including the widespread issue of fabricated experimental results, pointing to key challenges for future research.
@@ 2 MLR-Bench
Each step of the stepwise evaluation requires a set of input data for the agents. For example, in step (1), the Idea Generation agents require research tasks as input data; we use the same 201 tasks used in end-to-end evaluation. In step (2) the Proposal Generation agents require tasks and research ideas as input. To construct this dataset, for each of the 201 tasks, we randomly sample one idea (generated by any Idea Generation agent) from step (1) for this task, resulting in 201 (task, idea) pairs that serve as the input for step (2). We apply the same procedure in steps (3) and (4), each time randomly sampling from the outputs of the previous step for a task, except in step (3) we manually select a subset of 10 (task, idea, proposal) triples based on their suitability for evaluating the experimentation agents, to keep the experiments manageable. Details about these 10 selected tasks can be found in Section A.1.
@@ 2.1 MLR-Judge
Additionally, MLR-Judge incorporates an LLM judge alongside our rubrics. In this work, we select Gemini-2.5-Pro-Preview and Claude-3.7-Sonnet as our judge models, which have demonstrated superior reasoning capabilities and multimodal functionality, enabling them to review papers containing figures. These two judge models independently evaluate the quality of research outputs, and the final evaluation results are averaged from their individual assessments. At each research step, MLR-Judge evaluates the Consistency dimension by analyzing both the history information and the output of MLR-Agent. For the experimentation step, we examine the execution log of our coding agent to gain insights into the experimental process. Details of rubrics and prompts of MLR-Judge can be found in Section B.2.
@@ 2.2 MLR-Agent
MLR-Agent operates in two modes: a stepwise execution mode and an end-to-end execution mode. The stepwise execution mode enables us to evaluate each model’s performance in individual research steps, while the end-to-end execution mode allows us to assess the models’ capabilities in automating complete research projects. MLR-Agent’s automated research process follows the same research steps as illustrated in Fig. 1, namely (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing. MLR-Agent uses LLMs for steps (1) and (2), a coding agent (Claude Code) for step (3), and multimodal LLMs for step (4). Additionally, since most frontier models currently lack web search capabilities, they face challenges in accessing recent related work. Therefore, MLR-Agent uniformly employs GPT-4o-Search-Preview (searchgpt2024, ) to conduct a literature review between (1) Idea Generation and (2) Proposal Generation, providing reference materials for the agent. Details about the prompt used for each stage can be found in Section B.1.
@@ 3 How Well Can AI Agents Conduct Open-Ended Research?
This section describes empirical studies of how well AI agents perform on open-ended research. We use MLR-Agent as our agent scaffold to conduct research and evaluate a number of frontier models, as shown in Fig. 3. In experimentation and analysis step, our coding agent executes on an Ubuntu 22.04 server with access to four NVIDIA RTX 3090 GPUs. We use MLR-Judge to evaluate the outcomes of different research processes. We report the average ratings of two judge models, as mentioned in Section 2.1, and details of each rating can be found in Section A.3.
@@ Appendix A Experimental Details
A.1 Description of 10 Selected Tasks
@@
A.2 Comparison of AI Scientist V2 and MLR-Agent
We evaluate AI Scientist V2 Yamada2025ai , another agent scaffold with the capability of workshop-level end-to-end automated research, on MLR-Bench and compare its performance with MLR-Agent. We use o4-mini as the backbone model in the end-to-end automated research of AI Scientist V2. For MLR-Agent, we use o4-mini in idea generation, proposal generation and paper writing, and leverage Claude Code powered by claude-3.7-sonnet in experimentation and analysis. In this experiment, we only select five tasks iclr2025_bi_align, iclr2025_buildingtrust, iclr2025_dl4c, iclr2025_scsl and iclr2025_verifai from Table 7 and in the end-to-end evaluation, we does not attach any supplementary code along with the paper.
Table 8: End-to-End Performance Comparison of AI Scientist V2 and MLR-Agent. Each score is averaged across 5 tasks.
@@ A.2 Comparison of AI Scientist V2 and MLR-Agent
Gemini-2.5-Pro-Preview Judge
AI Scientist V2 $6.80\pm 1.17$ $6.80\pm 0.75$ $4.60\pm 1.20$ $6.20\pm 1.17$ $5.00\pm 0.89$
MLR-Agent $5.80\pm 0.98$ $6.80\pm 0.75$ $3.60\pm 1.20$ $5.20\pm 1.72$ $3.80\pm 1.47$
Claude-3.7-Sonnet Judge
AI Scientist V2 $7.00\pm 0.00$ $6.40\pm 0.80$ $5.60\pm 0.80$ $5.40\pm 1.02$ $5.60\pm 0.80$
MLR-Agent $8.00\pm 0.00$ $7.00\pm 0.00$ $6.00\pm 0.00$ $7.00\pm 0.00$ $6.80\pm 0.40$
Average of Gemini Judge & Claude Judge
AI Scientist V2 $6.90\pm 0.59$ $6.60\pm 0.55$ $5.10\pm 0.72$ $5.80\pm 0.78$ $5.30\pm 0.60$
MLR-Agent $6.90\pm 0.49$ $6.90\pm 0.38$ $4.80\pm 0.60$ $6.10\pm 0.86$ $5.30\pm 0.76$
We report the evaluation results of MLR-Judge powered by gemini-2.5-pro-preview and claude-3.7-sonnet and their average scores in Table 8. The average cost of o4-mini across these 5 tasks is 1.73 USD in AI Scientist V2 and 1.00 USD in MLR-Agent. From this table, we observe that Gemini Judge favors AI Scientist V2 across almost all review dimensions while Claude Judge shows the opposite preference. When averaged across both judge models, the two scaffolds demonstrate the same overall performance. The experimental results indicate that MLR-Agent offers comparable performance to AI Scientist V2 at a lower cost.
A.3 Evaluation Results of Different Judge Models
@@
This section shows the evaluation scores of different models in idea generation (see Tables 9 and 10), proposal generation (see Tables 11 and 12), paper writing (see Tables 13 and 14) and end-to-end evaluation (see Tables 15 and 16). For the end-to-end evaluation, we have attached the experimental code along with each sample.
@@
Table 13: Evaluation results judged by Gemini-2.5-Pro-Preview in paper writing.
@@
Table 15: End-to-end evaluation results judged by Gemini-2.5-Pro-Preview.
@@
Appendix B Prompt
@@ 1. Title: <title> (<arxiv_id>)
B.2 Prompts and Rubrics for MLR-Judge
@@ 6. Confidence Score (1-5)
B.2.2 Evaluation Prompts for MLR-Judge
@@
Appendix C Human Study Details

### v2

@@ 1 Introduction
  • Comprehensive Evaluation and Insights: We apply MLR-Bench to evaluate 6 frontier models, providing comparative insights into their strengths and limitations in scientific discovery. Also, we compare MLR-Agent with AI Scientist V2 on MLR-Bench, and find existing research agents exhibit critical failure modes, including the widespread issue of fabricated experimental results and hallucinated methodology, pointing to key challenges for future research.
@@ 2 MLR-Bench
Each step of the stepwise evaluation requires a set of input data for the agents. For example, in step (1), the Idea Generation agents require research tasks as input data; we use the same 201 tasks used in end-to-end evaluation. In step (2) the Proposal Generation agents require tasks and research ideas as input. To construct this dataset, for each of the 201 tasks, we randomly sample one idea (generated by any Idea Generation agent) from step (1) for this task, resulting in 201 (task, idea) pairs that serve as the input for step (2). We apply the same procedure in steps (3) and (4), each time randomly sampling from the outputs of the previous step for a task, except in step (3) we manually select a subset of 10 (task, idea, proposal) triples based on their suitability for evaluating the experimentation agents, to keep the experiments manageable. Details about these 10 selected tasks can be found in Appendix A.
@@ 2.1 MLR-Judge
Additionally, MLR-Judge incorporates an LLM judge alongside our rubrics. In this work, we select Gemini-2.5-Pro-Preview and Claude-3.7-Sonnet as our judge models, which have demonstrated superior reasoning capabilities and multimodal functionality, enabling them to review papers containing figures. These two judge models independently evaluate the quality of research outputs, and the final evaluation results are averaged from their individual assessments. At each research step, MLR-Judge evaluates the Consistency dimension by analyzing both the history information and the output of MLR-Agent. For the experimentation step, we examine the execution log of our coding agent to gain insights into the experimental process. Details of rubrics and prompts of MLR-Judge can be found in Section E.2.
@@ 2.2 MLR-Agent
MLR-Agent operates in two modes: a stepwise execution mode and an end-to-end execution mode. The stepwise execution mode enables us to evaluate each model’s performance in individual research steps, while the end-to-end execution mode allows us to assess the models’ capabilities in automating complete research projects. MLR-Agent’s automated research process follows the same research steps as illustrated in Fig. 1, namely (1) Idea Generation, (2) Proposal Generation, (3) Experimentation, and (4) Paper Writing. MLR-Agent uses LLMs for steps (1) and (2), a coding agent (Claude Code) for step (3), and multimodal LLMs for step (4). Additionally, since most frontier models currently lack web search capabilities, they face challenges in accessing recent related work. Therefore, MLR-Agent uniformly employs GPT-4o-Search-Preview (searchgpt2024, ) to conduct a literature review between (1) Idea Generation and (2) Proposal Generation, providing reference materials for the agent. Details about the prompt used for each stage can be found in Section E.1.
@@ 3 How Well Can AI Agents Conduct Open-Ended Research?
This section describes empirical studies of how well AI agents perform on open-ended research. We use MLR-Agent as our agent scaffold to conduct research and evaluate a number of frontier models, as shown in Fig. 3. In experimentation and analysis step, our coding agent executes on an Ubuntu 22.04 server with access to four NVIDIA RTX 3090 GPUs. We use MLR-Judge to evaluate the outcomes of different research processes. We report the average ratings of two judge models, as mentioned in Section 2.1, and details of each rating can be found in Appendix D.
@@ Appendix A Experimental Details
Description of 10 Selected Tasks
@@
Appendix B Comparison of AI Scientist V2 and MLR-Agent
In this section, we evaluate AI Scientist V2 Yamada2025ai , an agent scaffold designed for workshop-level end-to-end automated research, on the ten tasks listed in Table 7, and compare its performance with that of MLR-Agent. For AI Scientist V2, we used o4-mini as the backbone model for the whole pipeline and for each task, only a single attempt was made to generate the paper. For MLR-Agent, we also used o4-mini for idea generation, proposal generation, and paper writing, while leveraging Claude Code—powered by Claude-3.7-Sonnet—for experimentation and analysis.
Table 8: End-to-end performance comparison of AI Scientist V2 and MLR-Agent. Each score is averaged across 10 tasks. We report the average ratings of Gemini judge and Claude judge. The average costs across these 10 tasks are 1.88 USD for AI Scientist V2 and 1.15 USD for MLR-Agent.
@@ Appendix B Comparison of AI Scientist V2 and MLR-Agent
AI Scientist V2 $6.55\pm 0.94$ $6.70\pm 0.48$ $3.70\pm 1.29$ $4.85\pm 1.08$ $4.25\pm 1.25$
MLR-Agent $7.65\pm 0.32$ $6.90\pm 0.30$ $3.80\pm 0.60$ $5.05\pm 0.57$ $3.95\pm 0.71$
Main Results.
Table 8 summarizes the evaluation results from the Gemini and Claude judges. On average, AI Scientist V2 slightly outperforms MLR-Agent, although preferences vary across judges—each judge tends to favor a different model. However, both agents receive average scores below our acceptance threshold (i.e., 6 out of 10). A detailed breakdown reveals that while the agents perform well on Clarity and Novelty, their scores for Soundness and Significance are consistently below the threshold. Despite these limitations in research quality, both agents offer cost-effective solutions for large-scale experimentation, with average costs across the 10 tasks of \$1.88 for AI Scientist V2 and \$1.15 for MLR-Agent.
Key Findings.
These results lead to three key findings:
  • The agents excel at both generating novel and original ideas and presenting them clearly. Both AI Scientist V2 and MLR-Agent demonstrate a capacity for creativity, reflected in their high scores for Novelty. This indicates that they can propose original research questions and ideas crucial for scientific discovery. Also, the agents’ strong performance in the Clarity dimension shows they can articulate these ideas in a well-organized and easily understandable manner;
  • The methods and experiments proposed by the agents lack technical soundness. A primary weakness of AI Scientist V2 and MLR-Agent is that the proposed methods or techniques often lack soundness, which undermines the overall reliability of their research. This finding comes from their markedly low scores in the Soundness dimension, which consistently fall below the acceptance threshold. It suggests that while the agents can formulate creative ideas, the real implementations are frequently flawed;
  • The research conducted by the agents demonstrates limited potential impact. This limitation is primarily reflected in their consistently low scores for Significance, suggesting the work fails to address important research questions. This lack of perceived contribution, potentially compounded by difficulties in reproducing the experimental results, severely restricts the research’s influence and potential uptake by the scientific community.
Appendix C Hallucination Analysis of Existing Research Agents
To further investigate the issues in the content generated by AI Scientist V2 and MLR-Agent, we conducted a hallucination analysis in this section. Specifically, we aim to examine how frequently these research agents (i.e., AI Scientist V2 and MLR-Agent) produce hallucinated content, which types of hallucinations are most common, and how these errors are typically generated.
We observe that AI-generated research papers can fail to be trustworthy in several ways: for example, when the proposed method fails to address the stated motivation, or when the conclusions do not logically follow from the methodology. Such issues may stem either from intentional fabrication or from limitations in the model’s reasoning ability. To focus on objectively verifiable errors, we propose four frequently observed, fact-based hallucination types:
  • Faked Experimental Results: The data, metrics, or experiments’ outcomes are fabricated or never actually performed.
  • Hallucinated Methodology: The technical approaches are proposed but are not implemented (e.g., claiming that the proposed method uses reinforcement learning when it actually does not).
  • Incorrect Citations: The references to academic papers are incorrect or cannot be found.
  • Mathematical Errors: Incorrect equations, flawed derivations, or improper applications of mathematical concepts.
Figure 6: Frequency of each hallucination type in AI Scientist V2 and MLR-Agent across 10 tasks.
Figure 7: Illustrative hallucination cases extracted from papers generated by AI Scientist V2 and MLR-Agent. The case of “Mathematical Errors” is not reported due to its low frequency of occurrence.
To efficiently analyze the hallucination content, we first leverage two advanced automated judges—Gemini (Gemini-2.5-Pro-Preview-06-05) and Claude (Claude-3-7-Sonnet-20250219) to identify whether each generated paper contains any of these four hallucination types. Then the judgments are verified by our human annotators. To reduce the annotation burden, human annotators are provided with two review files, one from the Gemini judge and one from the Claude judge, which contain AI-detected hallucination types and their corresponding evidence. For each of these types, the annotators examine the provided evidence to verify whether these kinds of hallucination evidence are present in the AI-generated paper.
Key Observations.
Fig. 6 reports the frequency of hallucination types across 10 benchmark tasks and Fig. 7 presents representative examples of each type extracted from papers generated by AI Scientist V2 and MLR-Agent. We have two key observations.
Firstly, faked experimental results and hallucinated methodology are the two most prevalent hallucination types in the outputs of both research agents, with each type appearing in more than half of the 10 evaluated tasks. Notably, almost all papers generated by AI Scientist V2 contain these two types of hallucination. As illustrated in Fig. 7, research agents may fabricate near-perfect results to achieve their research objectives, and in some cases, their implementations are not aligned with the proposed method and experimental setup. One possible reason is that the generated ideas lack feasibility, leaving the coding agents unable to utilize available resources to achieve the experimental objectives. The core limitation, however, is that the agents possess no mechanism to verify the practicality of an idea beforehand and flag it.
Secondly, nonexistent citations are a significant issue, appearing in 50% of the tasks completed by MLR-Agent. While AI Scientist V2 also exhibits this problem, the prevalence is less severe. We hypothesize that this discrepancy stems from limitations in MLR-Agent’s literature collection tool, which restricts its ability to conduct accurate searches for relevant literature on the web. Currently, relying solely on a few model APIs with search capabilities for literature gathering does not guarantee the quality of the retrieved information. Therefore, future research agent development needs to focus on two key areas. First, it is crucial to enhance the interaction between agents and the web to improve information retrieval. Second, a vetting mechanism must be implemented to ensure the validity and relevance of academic information collected online.
Appendix D Evaluation Results of Different Judge Models
@@
This section shows the evaluation scores of different models judged by Gemini and Claude models. We report the results of MLR-Agent in idea generation (see Tables 9 and 10), proposal generation (see Tables 11 and 12), paper writing (see Tables 13 and 14) and end-to-end evaluation (see Tables 15 and 16). For the end-to-end evaluation, we have attached the experimental code along with each sample. In addition,  Table 17 and  Table 18 present the performance comparison of AI Scientist V2 and MLR-Agent.
@@
Table 13: Evaluation results judged by Gemini-2.5-Pro-Preview-05-06 in paper writing.
@@
Table 15: End-to-end evaluation results judged by Gemini-2.5-Pro-Preview-05-06.
@@
Table 17: End-to-end performance comparison of AI Scientist V2 and MLR-Agent judged by Gemini-2.5-Pro-Preview-05-06. Each score is averaged across 10 tasks.
Review Dimensions
Agent Scaffold Clarity Novelty Soundness Significance Overall
AI Scientist V2 $6.50\pm 1.69$ $6.90\pm 0.83$ $5.00\pm 2.37$ $5.50\pm 1.75$ $5.00\pm 2.19$
MLR-Agent $7.30\pm 0.64$ $6.80\pm 0.60$ $2.00\pm 1.00$ $3.50\pm 0.92$ $2.20\pm 1.17$
Table 18: End-to-end performance comparison of AI Scientist V2 and MLR-Agent judged by Claude-3.7-Sonnet-20250219. Each score is averaged across 10 tasks.
Review Dimensions
Agent Scaffold Clarity Novelty Soundness Significance Overall
AI Scientist V2 $6.60\pm 0.80$ $6.50\pm 0.50$ $2.40\pm 1.02$ $4.20\pm 1.25$ $3.50\pm 1.20$
MLR-Agent $8.00\pm 0.00$ $7.00\pm 0.00$ $5.60\pm 0.66$ $6.60\pm 0.66$ $5.70\pm 0.78$
Appendix E Prompt
@@ 1. Title: <title> (<arxiv_id>)
E.2 Prompts and Rubrics for MLR-Judge
@@ 6. Confidence Score (1-5)
E.2.2 Evaluation Prompts for MLR-Judge
@@
Appendix F Human Study Details